The speed of a car is decreased uniformly from 30 meters per second to 10 meters per second in 4 seconds. The magnitude of the car's acceleration in m/s^2 is


1
Expert's answer
2022-03-30T13:50:52-0400

v=u+at

Acceleration

"a=\frac{v-u}{t}"

"a=\frac{10-30}{4}=\frac{-20}{4}=-5m/sec^2"

Deaccleration